You have 40 individual socks in your sock drawer—10 black, 12 brown, 16 white, and 2 blue.
What is the probability of getting either a blue pair or a black pair in two tries?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52848)
Answer by greenestamps(13203)



The statement of the problem does not define what "in two tries" means -- leaving the problem open to different interpretations.

I read the problem differently than the other tutor; my interpretation is that "in two tries" means you are selecting two of the 40 individual socks from the drawer, one after the other, without replacement.

P(blue pair) = (2/40)(1/39) = 2/1560
P(black pair) = (10/40)(9/39) = 90/1560
P(blue pair or black pair) = 2/1560+90/1560 = 92/1560 = 23/390

ANSWER: 23/390 or about 5.9%